Assume a large primary system that consist of a vessel containing warm water in the system with a temperature of 350,001k. An infinitesimal amount of its heat of 1050j is transferred to a secondary system consisting of a room air at a temperature of 350k. Consequently the water temperature drops to 349,999k. What is the entropy generation: A) If the system is cooled in air. B) If the heat removed from the primary system is used to run a cannot engine process and that rejected by the engine is transferred to the secondary system.
1
Expert's answer
2020-04-27T01:55:19-0400

(A) Air temperature is at 350 K, Heat transfer= dQ= 1050 J

Entropy generation,

"S_{gen}= \Delta S_{sys}+ \frac{dQ}{T}"

Here, in case of ambient condititon change in entropy will be zero

"S_{gen}=\frac{1050}{350}=3 \frac{J}{K}"

(B) As per the second case we know that for carnot heat engine

"\frac{Q_1}{Q_2}= \frac{T_1}{T2}"


"\frac{1050}{Q_2}= \frac{350.001}{349.999}"


Q2=1049.99 J

and entropy generation

"S_{gen}=\frac{1049.999}{349.999}= 2.999 J/K"